[Brahms-l] Shift report 20040102 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Fri Jan 02 2004 - 08:11:40 EST
------------------------------------------------------------
Date: Fri Jan  2 08:11:37 2004, Shift: 00:00-08:00
Supervisor: Catalin Ristea, with: 
Log book#15 pages 20 - 20

Shift Summary:
------------------------------------------------------------
1. Long store running from the previous shift.
2. Run 9312 - D1=D2=D3=D4=1/4B, D5=350B - 280k(tr1) events
   run 9313 - D1=D2=D3=D4=1/3B, D5=350B - 315k(tr1) events
   run 9314 - D1=D2=D3=D4=1/3B, D5=350B -  50k(tr1) events
   run 9315 - D1=D2=D3=D4=1/3B, D5 no field - 46k(tr1) ev.
   run 9316 - D1=D2=D3=D4=1/3B, D5=1050B - running
3. No new problems with the detectors.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jan 2 08:12:03 2004

This archive was generated by hypermail 2.1.8 : Fri Jan 02 2004 - 08:12:17 EST